# Learning Visual Hierarchies in Hyperbolic Space for Image Retrieval

### v2

@@ Abstract
Figure 1: An illustration of part-based image hierarchy organized in hyperbolic space. At the highest level, we see the urban environment, composed of buildings, streets, and sky. Zooming in, we find the building category, which further divides into skyscrapers, mid-rise structures, and more. Each of them has its own visual elements, which in turn can be decomposed into sub-elements. Best viewed when zoomed in.
@@ 1 Introduction
For experimentation, we build a dataset of visual hierarchies using the bounding box annotations of OpenImages [28]. Our dataset includes entailment relationships between scenes, objects, and parts, within a single image as well as across images at the category level. Similarly, for hierarchical retrieval evaluation, we use the full training set to create ground truth hierarchy trees per scene/object. Additionally, we design a metric for evaluating hierarchical retrieval based on the optimal transport distance between the label distribution of the retrieval set and ground truth label distribution within the hierarchy tree. Combined with Recall@k metrics, this demonstrates that our method captures semantic and structural information, transcending mere visual similarity. Furthermore, to the best of our knowledge, our model is the first to generalize to out-of-domain image hierarchies, achieving strong performance on unseen and diverse datasets.
@@ 2 Preliminaries
$\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{H}}=-x_{0}y_{0}+\sum_{i=1}^{d}x_{i}y_{i}\ .\vspace{-1ex}$ (2)
Here, the 00-th dimension of the vector is treated as the time component and the rest as the space component. From the definition of $\mathbb{H}^{d}$, the time component can be written using the space component as follows:
@@
[…]
@@
Figure 2: An illustrative example image hierarchies. a) Image $I$ with object-level bounding boxes. Each bounding box is entailed by $I$. b) Hierarchies created via bounding box-to-bounding box entailment within $I$ (larger bounding boxes entail smaller ones). c) Cross-image hierarchy created by sampling $N$ bounding boxes with corresponding object classes from other images, which are then entailed by $I$. Find details of cross-image sampling in Sec. 3.1.
@@ 3.1 Part-Based Image Hierarchy
Our entailment rule above naturally facilitates a pairwise relationship. Let $I\in\mathcal{I}$ be an image, either the full scene image or a cropped bounding box, and let $\mathcal{B}$ denote the set of all bounding boxes in the dataset. Suppose $\mathcal{B}_{I}$ be the set of bounding boxes contained in the image $I$, i.e, $\mathcal{B}_{I}=\{b\in\mathcal{B}\mid b\subset I\}$ [^1] We use the $\subset$ notation to denote contained in relationship. In the case where $I$ is a cropped bounding box, this relationship is defined to hold if the majority (e.g., 80%) of the small bounding box $b$ is contained in $I$.. Note, each bounding box has an associated object label denoted by $b_{l}\in\mathcal{L}$.
@@ 3.2 Angle-Based Entailment Loss
$-\mathbb{E}_{\mathcal{D}}\!\left[\log\frac{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)}{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)+\sum_{\mathbf{y}^{-}\in\mathcal{N}_{i}}\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i}^{-})}{\tau}\right)}\right]$
@@ 3.3 Hierarchical Retrieval Evaluation
where ¯$\bar{\cdot}\in\mathbf{R}^{m+1}$ is the label distribution with other class added [^2] For $\mathbf{h}_{I}$ other class has zero mass, and for $\mathbf{r}_{I}$ all labels not in $\mathcal{H}_{I}$ are combined to form the other class.. Note that a smaller distance indicates better alignment.
@@ 5 Experiments
For training and evaluation, we construct HierOpenImages, a novel dataset containing pairwise part-based image hierarchies built from the OpenImages dataset [28]. We further evaluate generalization on out-of-domain unseen datasets and hierarchies on the LVIS [17] dataset and 10 popular single-class datasets.
@@ 5.1 Main Results
[…]
@@
Table 2: Part-based hierarchical evaluation of parent-to-child image retrieval on HierOpenImages. Results are evaluated using the ground truth hierarchy tree and the hierarchical distribution of the test set. Smaller OT distance indicates better distribution alignment (ref. Sec. 3.3).
@@
Fig. 5 shows the Precision-Recall (PR) curve with increasing $\beta_{1}$ angle thresholds (00 to $\pi$) for CLIP-hyp and CLIP-euc, and with cosine similarity thresholds (0 to 1) for the baseline CLIP. The proposed hyperbolic model trained without cross-image samples substantially outperforms the Euclidean model and even exceeds the Euclidean model with cross-image samples, indicating it implicitly learns the underlying structure. When cross-image samples are introduced, the hyperbolic model still leads, though the performance gap narrows.
Figure 4: Example of parent-to-child retrieval using CLIP ViT and our CLIP-hyp model. Results are ordered by ascending norms. Our model retrieves images matching the predefined scene-object-part hierarchy, placing high-level objects near the origin (e.g, harbor $\rightarrow$ boart parts), and grouping semantically related but visually distinct objects (e.g, microwave oven & kitchen hood).
@@
Following [9, 15, 40], we use parent-to-child image traversals with results ordered by increasing embedding norm, to illustrate the latent space, which: (1) aligns with the predefined scene-object-part hierarchy, placing high-level objects near the origin. (2). groups diverse objects under the same lower hierarchical branch, even if they are visually distinct (e.g., keyboard and monitor under the studio image). In contrast, CLIP struggles to link large scenes with smaller objects.
@@ 5.2 Generalization Evaluation
We evaluate the generalization of our method through image retrieval on unseen datasets and hierarchies. Similar to Table 1-2, Table 3 presents part-based same-class and hierarchical retrieval performance on the out-of-domain LVIS dataset [17], which contains 1,203 fine-grained object classes, including many rare objects absent from the training HierOpenImages dataset. Table 4 evaluates image retrieval performance on 10 popular single-class datasets using top-5 majority voting. Results demonstrate that our visual hierarchical learning significantly enhances model generalization on unseen datasets and image hierarchies.
@@
Model
CIFAR-10 [27]
CIFAR-100 [27]
SUN397 [49]
Caltech-101 [29]
STL-10 [11]
Tiny-ImageNet [8]
Food101 [3]
CUB [47]
Pets [37]
Flowers [34]
CLIP 55.1 28.7 39.7 62.2 63.5 11.8 15.0 14.3 31.6 40.0
CLIP-euc 80.2 44.8 31.8 78.3 93.1 29.2 23.8 4.8 11.6 7.0
CLIP-hyp 84.1 49.7 38.3 79.7 94.7 33.3 42.3 20.5 34.4 36.9
Table 4: Unseen single-class image retrieval (image-only model). For each query image, retrieve the top-5 images based on cosine similarity, and predict the class by majority vote among these top images.
@@ Appendix A Hierarchy Tree
Our part-based image hierarchy framework is designed to be widely applicable to general image datasets with bounding box annotations. In this section, we outline the key implementation details involved in constructing hierarchy trees.
We introduce a general method to generate dataset-specific ground truth hierarchy trees based on data statistics. Following the approach outlined in Sec. 3.1 of the main paper, we begin by identifying bounding box pairs with substantial overlap. In this work, we define a bounding box pair when at least 80% of the smaller bounding box $b$ is contained within either the full image or a larger bounding box. We initially set the overlap threshold at 100% and empirically adjusted it by evaluating the hierarchy’s validity using text labels. For the OpenImages dataset, 80% was found to be the most suitable threshold. This is a design choice and can be adapted for other datasets.
@@ B.2 Part-based Image Retrieval
The HierOpenImages test set contains approximately 1 million full images and 3 million bounding box images. During part-based image retrieval evaluation, we filter out bounding boxes for very small or large objects that occupy less than 5% or more than 30% of the full image area. This reduces the candidate full images to 59,691 and bounding box images to 330,063. For part-based image retrieval, we randomly select a subset of 10,000 full images and 10,000 bounding box images as query images, using the entire test set as candidates. The bounding box images in the query set span 339 object classes, while the full images are labeled with 516 classes. The top-50 class frequency distributions for both the query and candidate sets are shown in Fig. 7. Although the query and candidate set distributions are similar, the class distribution is highly imbalanced, highlighting the importance of hierarchical retrieval evaluation using combined precision-recall curves and OT distances (Sec. 3.3; see Fig. 5 and Table 2 in the main paper).
@@
Model
CIFAR-10 [27]
CIFAR-100 [27]
SUN397 [49]
Caltech-101 [29]
STL-10 [11]
Tiny-ImageNet [8]
Food101 [3]
CUB [47]
Pets [37]
Flowers [34]
Query N 50,000 50,000 1,829 1,829 8,000 10,000 25,250 5,794 3,669 6,149
Candidate N 10,000 10,000 7,315 7,315 5,000 100,000 75,750 5,994 3,680 1,020
Table 5: Query and candidate set sizes for each single-class unseen dataset.
@@
In Table 2 of the main paper, we retrieve the Top-K results starting from approximately $50\%$ of the test data size. This is necessary because each image contains an average of 5.29 distinct classes and 61.5 classes across hierarchical trees, requiring a large number of retrievals to accurately evaluate the distribution.
Figure 8: Example of our image retrieval interface built with Gradio [1]. The interface supports image selection/upload and retrieves results ranked by user-defined modes. Its modular design allows for easy integration of additional functionalities.
@@ B.3 Generalization Evaluation
We evaluate the generalization capability of our model on the out-of-domain LVIS dataset [17], which is designed for long-tail instance classification and segmentation. It has a highly imbalanced distribution of 1,203 object categories and contains 897 object categories that are absent from OpenImages [28]. The full list of these categories can be found in the appendix. Only display the first class of synonyms.
We construct the hierarchical evaluation set from the validation set of the LVIS dataset [17], which contains 19,809 images. Following the same process as constructing HierOpenImages, we evaluate bounding box images and corresponding full images that occupy 5%-30% of the full image area, reducing the candidate set to 14,716 full images and 76,255 bounding box images. We use the full reduced set for part-based image retrieval in Table 3 in the main paper.
[…]
Single-class Datasets.
We use the official splits from torchvision, designating the training set as candidates and the test/validation set as queries. For datasets without predefined splits on torchvision, such as SUN397 [49] and Caltech101 [11], we randomly split the dataset into $80\%$ candidates and $20\%$ queries. The sizes of query and candidate images for each dataset are listed in Table. 5.
For each dataset, we retrieve the top-5 most similar images from the candidate set based on cosine similarity and predict the class by majority vote among top images. We chose cosine similarity instead of hyperbolic angles as our metric, as there is no part-based relationship between the query and candidate images, and this is not an evaluation of preserving a pre-defined hierarchy.
Appendix C Ablation Studies
In this supplementary material, we further evaluate these models on part-based same-class and hierarchical image retrieval tasks, as shown in Table. 6-7. The best results are highlighted in bold and the second-best results are shown in blue. The results clearly indicate that cross-image sampling improves image retrieval performance. Notably, the hierarchical retrieval results align with Fig. 5 in the main paper. Specifically, the hyperbolic model trained without cross-image samples outperforms the Euclidean model trained with cross-image samples in 5 out of 6 cases, as shown in Table 7.
@@ Appendix D More Qualitative Results
Figure 9: Example of parent-to-child retrieval using CLIP ViT and our CLIP-hyp model. Results are ordered by ascending embedding norms. Our model retrieves images matching the predefined scene-object-part hierarchy, placing high-level objects near the origin (e.g, group of fruits $\rightarrow$ single fruits), and grouping semantically related but visually distinct objects (e.g, chairs & TVs). All retrieved bounding box images are scaled to the same ratio.

### v3

@@ Abstract
work done while at Amazon
Figure 1: An illustration of part-based image hierarchy organized in hyperbolic space. At the highest level, we see the urban environment, composed of buildings, streets, and sky. Zooming in, we find the building category, which further divides into skyscrapers, mid-rise structures, and more. Each of them has its own visual elements, which in turn can be decomposed into sub-elements.
@@ 1 Introduction
For experimentation, we build a dataset of visual hierarchies using the bounding box annotations of OpenImages [24]. Our dataset includes entailment relationships between scenes, objects, and parts, within a single image as well as across images at the category level. Similarly, for hierarchical retrieval evaluation, we use the full training set to create ground truth hierarchy trees per scene/object. Additionally, we design a metric for evaluating hierarchical retrieval based on the optimal transport distance between the label distribution of the retrieval set and ground truth label distribution within the hierarchy tree. Combined with Recall@k metrics, this demonstrates that our method captures semantic and structural information, transcending mere visual similarity. Furthermore, to the best of our knowledge, our model is the first to generalize to out-of-domain image hierarchies, achieving strong performance on unseen and out-of-domain datasets.
@@ 2 Preliminaries
$\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{H}}=-x_{0}y_{0}+\sum_{i=1}^{d}x_{i}y_{i}\ .\vskip-4.30554pt$ (2)
Here, the 0-th dimension of the vector is treated as the time component and the rest as the space component. From the definition of $\mathbb{H}^{d}$, the time component can be written using the space component as follows:
@@
where $\|\cdot\|$ is the Euclidean norm and $\mathbf{x}_{\text{space}}=\mathbf{x}_{1:d}$}.
@@
Figure 2: An illustrative example image hierarchies. a) Image II with object-level bounding boxes. Each bounding box is entailed by $I$. b) Hierarchies created via bounding box-to-bounding box entailment within II (larger bounding boxes entail smaller ones). c) Cross-image hierarchy created by sampling NN bounding boxes with corresponding object classes from other images, which are then entailed by $I$. Find details of cross-image sampling in Sec. 3.1.
@@ 3.1 Part-Based Image Hierarchy
Our entailment rule above naturally facilitates a pairwise relationship. Let $I\in\mathcal{I}$ be an image, either the full scene image or a cropped bounding box, and let $\mathcal{B}$ denote the set of all bounding boxes in the dataset. Suppose $\mathcal{B}_{I}$ be the set of bounding boxes contained in the image $I$, i.e., $\mathcal{B}_{I}=\{b\in\mathcal{B}\mid b\subset I\}$ [^2] We use the $\subset$ notation to denote contained in relationship. In the case where $I$ is a cropped bounding box, this relationship is defined to hold if the majority (e.g., 80%) of the small bounding box $b$ is contained in $I$.. Note, each bounding box has an associated object label denoted by $b_{l}\in\mathcal{L}$.
@@ 3.2 Angle-Based Entailment Loss
$-\mathbb{E}_{\mathcal{D}}\!\left[\log\frac{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)}{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)+\sum_{\mathbf{y}^{-}\in\mathcal{N}_{i}}\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i}^{-})}{\tau}\right)}\right]\ ,$
@@ 3.3 Hierarchical Retrieval Evaluation
where $\bar{\cdot}\in\mathbf{R}^{m+1}$ is the label distribution with other class added [^3] For $\mathbf{h}_{I}$ other class has zero mass, and for $\mathbf{r}_{I}$ all labels not in $\mathcal{H}_{I}$ are combined to form the other class.. Note that a smaller distance indicates better alignment.
@@ 5 Experiments
For training and evaluation, we construct HierOpenImages, a novel dataset containing pairwise part-based image hierarchies built from the OpenImages dataset [24]. We further evaluate generalization on out-of-domain datasets and hierarchies.
@@ 5.1 Main Results
Denoting the full image as parent and the bounding box as child, we evaluate retrieval tasks in both child-to-parent and parent-to-child directions. Table 1 shows the retrieval accuracy of $\text{top-k}=\{5,10,50,100\}$. We notice a significant and consistent improvement with our proposed hyperbolic model, across all metrics and model variants. This highlights the relevance of our angle-based entailment loss and the advantages of learning hierarchical image embeddings in hyperbolic space. While for HCL [12], only a slight performance increase was observed after fine-tuning.
@@
Table 2: Part-based hierarchical evaluation of parent-to-child image retrieval on HierOpenImages. Results are evaluated using the ground truth hierarchy tree and the hierarchical distribution of the test set. Smaller OT distance indicates better distribution alignment.
@@
Fig. 4 shows the Precision-Recall (PR) curve with increasing $\beta_{1}$ angle thresholds (0 to $\pi$) for CLIP-hyp and CLIP-euc, and with cosine similarity thresholds (0 to 1) for the baseline CLIP. The proposed hyperbolic model trained without cross-image samples substantially outperforms the Euclidean model and even exceeds the Euclidean model with cross-image samples, indicating it implicitly learns the underlying structure. When cross-image samples are introduced, the hyperbolic model still leads, though the performance gap narrows.
Figure 5: Example of parent-to-child retrieval using CLIP ViT and our CLIP-hyp model. Results are ordered by ascending norms. Our model retrieves images matching the predefined scene-object-part hierarchy, placing high-level objects near the origin (e.g., harbor $\rightarrow$ boart parts), and grouping semantically related but visually distinct objects (e.g., microwave oven & kitchen hood).
@@
Following [7, 11, 33], we use parent-to-child image traversals with results ordered by increasing embedding norm, to illustrate the latent space, which: (1) aligns with the predefined scene-object-part hierarchy, placing high-level objects near the origin. (2). groups diverse objects under the same lower hierarchical branch, even if they are visually distinct (e.g., keyboard and monitor under the studio image). Moreover, Figure 10 shows an example of the model’s emergent structure where high-frequency, ambiguous image crops are naturally pushed toward the boundary of hyperbolic space. Find details in supplementary materials.
Figure 6: Example of the model’s emergent structure.
@@ 5.2 Generalization Evaluation
We evaluate the generalization of our method through image retrieval on unseen datasets and hierarchies. Similar to Table 1-2, Table 3 presents part-based same-class and hierarchical retrieval performance on the out-of-domain LVIS dataset [14], which contains 1,203 fine-grained object classes, including many rare objects absent from the training HierOpenImages dataset. Table 4 evaluates image retrieval performance on VOC and COCO datasets following the zero-shot object detection in [30] where ground-truth boxes are used as region proposals, and then predict via top-5 majority vote. Results demonstrate that our visual hierarchical learning significantly enhances model generalization on unseen datasets and image hierarchies.
@@
Dataset CLIP CLIP-euc CLIP-hyp
VOC 87.3 92.2 93.9
COCO 63.6 66.5 71.8
Table 4: Zero-shot object detection using ground-truth boxes. Image-only model with top-5 majority voting.
@@ Appendix A Hierarchy Tree
Our part-based image hierarchy framework is designed to be widely applicable to general image datasets with bounding box annotations. In this section, we outline the key implementation involved in constructing hierarchy trees.
We introduce a general method to generate dataset-specific ground truth hierarchy trees based on data statistics. Following the approach outlined in Sec. 3.1 of the main paper, we begin by identifying bounding box pairs with substantial overlap. In this work, we define a bounding box pair when at least 80% of the smaller bounding box $b$ is contained within either the full image or a larger bounding box. We initially set the overlap threshold at 100% and empirically adjusted it by evaluating the hierarchy’s validity using text labels. For the OpenImages dataset, 80% was found to be a suitable threshold. This is a design choice and can be adapted for other datasets.
@@ B.2 Part-based Image Retrieval
During part-based image retrieval evaluation, we filter out bounding boxes for very small or large objects, and filter out noisy samples (e.g., wrong labels, highly occluded, etc). For part-based image retrieval, we randomly select a subset of 10,000 full images and 10,000 bounding box images as query images, using the entire filtered test set as candidates. The top-50 class frequency distributions for both the query and candidate sets are shown in Fig. 8. Although the query and candidate set distributions are similar, the class distribution is highly imbalanced, highlighting the importance of hierarchical retrieval evaluation using combined precision-recall curves and OT distances (Sec. 3.3; see Fig. 4 and Table 2 in the main paper).
@@
In Table 2 of the main paper, we retrieve the Top-K results starting from a large K. This is necessary because each image contains an average of 5.29 distinct classes and 61.5 classes across hierarchical trees, requiring a large number of retrievals to accurately evaluate the distribution.
Figure 9: Example of our image retrieval interface built with Gradio [1]. The interface supports image selection, upload and retrieves results ranked by user-defined modes. Its modular design allows for easy integration of additional functionalities.
@@
Figure 10: Example of the model’s emergent structure. In the left plot, the percentage of samples of the original Sunglasses crops are shown in dark green. Applying Gaussian blur or selecting extremely small crops shifts the embeddings further from the root. The right figure illustrates a clear separation between full-image and part-image embeddings.
@@ B.3 Generalization Evaluation
We evaluate the generalization capability of our model on the out-of-domain LVIS dataset [14], which is designed for large-scale ($\sim$1.2M bounding boxes) long-tail instance classification and segmentation. It has a highly imbalanced distribution of 1,203 object categories and contains 897 object categories that are absent from OpenImages [24]. Here are some examples of unseen hierarchies in LVIS [14] but not in OpenImges [24]: {table $\rightarrow$ tablecloth $\rightarrow$ ashtray $\rightarrow$ cigarette}, {backpack $\rightarrow$ strap $\rightarrow$ belt buckle}, {toy $\rightarrow$ teddy bear $\rightarrow$ thread $\rightarrow$ bobbin}, {sofa $\rightarrow$ blanket $\rightarrow$ quilt $\rightarrow$ bedspread}. The full list of these categories can be found in the appendix. Only display the first class of synonyms.
We construct the hierarchical evaluation set from the validation set of the LVIS dataset [14], following the similar process as constructing HierOpenImages. To construct the ground truth hierarchy tree, we use the pipeline as described in Sec. A. We empirically choose parameter $\textit{frequency}=5$ and $\textit{proportion}=5\%$ in our experiments. This adjustment is necessary because the LVIS dataset [14] has very unbalanced classes.
Emergent structure.
We observe that high-frequency, ambiguous image crops are naturally pushed toward the boundary of hyperbolic space, where more angular entailment constraints can be better satisfied. This aligns with the exponentially growing volume near the boundary, and it is an efficient solution for representing diverse, overlapping semantics. In Figure 10, in the left plot, the percentage of samples of the original Sunglasses crops are shown in dark green. Adding Gaussian blur or selecting extremely small crops shifts their embeddings further from the root. The small crops (red curve) are not included in other curves. The right plot shows clear separation between full and part image embeddings.
@@
Appendix C Ablation Studies
In this supplementary material, we further evaluate these models on part-based same-class and hierarchical image retrieval tasks, as shown in Table. 5-6. The best results are highlighted in bold and the second-best results are shown in blue. The results clearly indicate that cross-image sampling improves image retrieval performance. Notably, the hierarchical retrieval results align with Fig. 4 in the main paper. Specifically, the hyperbolic model trained without cross-image samples outperforms the Euclidean model trained with cross-image samples in 5 out of 6 cases, as shown in Table 6.
@@ Appendix D More Qualitative Results
Figure 11: Example of parent-to-child retrieval using CLIP ViT and our CLIP-hyp model. Results are ordered by ascending embedding norms. Our model retrieves images matching the predefined scene-object-part hierarchy, placing high-level objects near the origin (e.g., group of fruits $\rightarrow$ single fruits), and grouping semantically related but visually distinct objects (e.g., chairs & TVs). All retrieved bounding box images are scaled to the same ratio.